The fifth-grade teachers took their classes on a field trip to a museum. The first group had 25 students and 2 teachers and cost $97.50. The sencond group had 32 students and 3 teachers and cost $127. Whats the cost oer student, per teacher?
Answer by macston(5194)

.
S=student cost; T=teacher cost
.
32S+3T=$127 . Multiply this by 2.
64S+6T=$254 Subtract this below
.
25S+2T=$97.50 . Multiply this by 3.
75S+6T=$292.50
64S+6T=$254 . Subtract this (from above.)
.
11S=$38.50
S=$3.50
ANSWER 1: Student tickets cost $3.50.
.
32S+3T=$127
32($3.50)+3T=$127
$112+3T=$127
3T=$15
T=$5.00
ANSWER 2: Teacher tickets cost $5.00.
.
CHECK:
25S+2T=$97.50
25($3.50)+2($5.00)=$97.50
$87.50+$10.00=$97.50
$97.50=$97.50